Outage: Connectivity issues – Brussels: Solved

Dear customers,

We are currently experiencing an outage with internet connectivity in Brussels region. Our Network Operations team is investigating this outage. Further updates will be released as soon as possible.

Please find all details below:

Start*: 27-01-2023 09:25
End*: 27-01-2023 16:11
Duration: 6h46min
Impacted services: All internet services
Affected Areas**: 02STR: Brussels

 

Update 11:53: We have identified the failure and are currently working with our supplier to resolve it as soon as possible. Our engineer is on the way to the location to replace the faulty card should the problem not be resolved remotely. More updates will follow soon.

Update 13:22: Our engineer has arrived to the location to swap the malfunctioning card. We will update you as soon as the works have been completed.

Update 14:20: The card has been replaced, but the affected sessions are still not able to connect. We are in contact with our supplier and are pushing them to pinpoint the core issue ASAP.

Update 15:55: The investigation is still ongoing. Our supplier just reported that they expect to have the sessions restored within 30 minutes – we are waiting for confirmation.

Update 16:19: The incident is resolved. Most sessions have already recovered, and traffic flow is coming back to normal. We are keeping an eye on the situation. If your modem has not been able to connect yet, a reboot may be needed.

Update 18:00: We noticed that some fiber customers in Etterbeek, Koekelberg and Molenbeek are still not able to connect. We will escalate this towards our supplier. More updates will follow.

We apologize for the inconveniences.

Best regards
The edpnet team

*All times are listed in CET, Central European Time
** Check the different areas & zones on our support site


Issue: High latency during peak hours – region: Antwerp, Limburg, Flemish Brabant: SOLVED

Dear customers,

We are currently experiencing an issue with internet services in Antwerp, Limburg, Flemish Brabant (telephone zones: 03, 011, 012, 013, 014, 015, 016, 089). Customers may notice service degradation and high latency during peak hours. Our Network Operations team is investigating the outage.

Update 15/11/2022:  We located the problem and our config changes appeared to have solved the problem, we will be keeping an eye on the situation.

Further details will be released as soon as possible.

We apologize for the inconveniences.

Best regards
The edpnet team


Outage: Internet & telephony in service area 3 down : SOLVED

Dear customers,

We are currently experiencing an outage with internet and telephony in service area 3.  Our Network Operations team is investigating this outage.  Further updates will be released as soon as possible.

Update 8:32: problem still under investigation, cause unclear, troubleshooting our equipment and contacting Proximus to do the same

Update 9:20: internet traffic is restoring back to normal. Connections should not be affected anymore. We are still monitoring in the background.

Update 10:40: we found a problem on our optical network, we brought down the links that were on that card and there is no impact since then. We will replace the card this afternoon.

Please find all details below:

Start*: 19-10-2022 02:30 (noticeable impact since around 7:00)
End*: 19-10-2022 08:45
Duration: 6 hours 15 minutes
Impacted services: Internet & telephony
Affected Areas**: Service area 3: 03, 011, 012, 013, 014, 015, 016, 089

We apologize for the inconveniences.

Best regards
The edpnet team

*All times are listed in CET, Central European Time
** Check the different areas & zones on our support site 


Outage : Mail delivery delayed : SOLVED

UPD 30/09/2022 08h53: The mail problem is resolved. Mail traffic is completely back to normal since 29/09/2022 18h00.

UPD 16h10: our supplier has identified the issue and has implemented a patch for mail gateways. The mail traffic is being restored but there are still delays of 2-3 hours.

UPD 12h12: the mail issue is a global problem at our supplier. Mail servers are affected all over the world. Our supplier is trying to fix the issue in the shortest possible time.

Dear customers,

We are currently experiencing problem with our mail servers due to a massive influx of spam.

Mail delivery is currently delayed. No mails should be lost.

Our Network operations team is working now on resolving the issue as soon as possible.

Start*: 29-09-2022 04:00
End*: 29-09-2022 18:00
Duration: 14 hours
Impacted services: mail services
Affected Areas**: all areas

We apologize for the inconveniences.

Best regards
The edpnet team

*All times are listed in CET, Central European Time
** Check the different areas & zones on our support site


Outage notification: possible issues with internet and voice services in West and East Flanders : SOLVED

UPD 14h25 : the fiber supplier has confirmed the end of the works. Traffic is running its normal path and no further interruptions are expected.

UPD 12h45 : the link has recovered and has been stable since 11:07. We are awaiting an official confirmation from the fiber supplier.

Dear customers,

We are currently experiencing a fiber cut in 91GEN area (West and East Flanders). Traffic is being rerouted but customers may notice minor issues with the internet and voice services.

Our Network Operations team is investigating the outage.

Further details will be released as soon as possible.

We apologize for the inconveniences.

Best regards
The edpnet team


Outage: Outbound voice issues (Solved)

Dear customers,

We experienced an outage with some of our outbound calls.  Our Network Operations team investigated and resolved this outage.

Please find all details below:

Start*: 04-04-2022 14:15*
End*: 04-04-2022 16:15*
Duration: 2hrs
Impacted services: Some outbound calls
Affected Areas**: All areas

We apologize for the inconveniences.

Best regards
The edpnet team

*All times are listed in CET, Central European Time
** Check the different areas & zones on our support site


Outage: Network hardware failure: Solved

Dear customers,

We are currently experiencing an outage with our network backbone.  Our Network Operations team is investigating this outage.  Further updates will be released as soon as possible.

Edit 18:50*:  Broken card replaced, all services restored.

Please find all details below:

Start*: 09-02-2022 13:49*
End*: 09-02-2022 18:31*
Duration: 5:18
Impacted services: All internet services
Affected Areas**: 02STR, 03BKC, 41GRE

 

Edit 16:15*: We have identified the failure and mitigated its downsides. Users may have noticed increased latency from 14:00 onwards but this should be resolved since 16:15. Our techs are moving onsite to replace affected hardware.
Edit 17:26*: You may still notice some latency, we are working on minimizing this as much as possible until the hardware replacement is done.
Edit 18:50*:  Broken card replaced, all services restored.

We apologize for the inconveniences.

Best regards
The edpnet team

*All times are listed in CET, Central European Time
** Check the different areas & zones on our support site


Outage: Mobile services: SOLVED

Dear customers,

We are currently experiencing an outage with our Mobile services.  Our network provider is investigating this outage.  Further updates will be released as soon as possible.

EDIT 19:06 Our provider solved the issue.

Please find all details below:

Start*: 19-11-2021 17:00
End*: 19-11-2021 19:01
Duration: 2h 01
Impacted services: Data Services, SMS, Voice Services
Affected Areas**: All areas

 

We apologize for the inconveniences.

Best regards
The edpnet team

*All times are listed in CET, Central European Time
** Check the different areas & zones on our support site


Outage: General DDoS on edpnet services

Dear customers,

We are currently experiencing an distributed denial of service attack (DDoS) on all of our services. This may cause you to experience problems with all of our services.  Our Network Operations team is investigating this outage.  Further updates will be released as soon as possible.

Please find all details below:

Start*: 15-09-2021 16:15
End*: 19-09-2021 22:24
Duration: //
Impacted services: All services are impacted
Affected Areas**: All areas

EDIT 16:44*: As soon as we identified the issue, we have started routing our data through a scrubbing service (which tries to identify harmful sources & data and removes it).

EDIT 17:19*: The situation seems to be improving as we identify more and more attack vectors and sources. It seems like for most impacted customers, the situation has improved markedly since approximately ~17h.

EDIT 16/09/2021 10:12*: During the night we had two more attacks. We are working with the authorities, who have confirmed they are looking into it and are doing everything in their power to find the responsible individuals. We were contacted by an individual who verified he was behind the attacks, asking for a ransom.

EDIT 13:18*: We are being DDoS’ed again. We are doing what we can to mitigate the issues.

EDIT 15:41*: The attacks are still on-going. To give you a perspective of what we are dealing with, this is what our filtering service has been able to filter out so far:

EDIT 16:39*: We are monitoring the situation closely, and are discussing the situation both with our anti-DDoS provider and Belgian official instances. We hope to improve the situation as soon as possible and will keep you in the loop of our progress.

EDIT 17/9 15:47*: The attacks have started again. All of our uplinks are entirely congested. We are trying to mitigate the DDoS once more.

EDIT 18/09 12:41*: Yesterday evening and past night we were hit with multiple attacks.  Currently no attacks since 03:00. We still expect more attacks to come.  We are monitoring 24/7.  And we will communicate accordingly.  We try to answer questions or comments on social media as good as we can.

EDIT 20/09 9:48: There were some attacks on Sunday afternoon. No new attacks since then. We keep monitoring and do our best to minimize impact in case of new spikes.

We apologize for the inconveniences.

Best regards
The edpnet team

*All times are listed in CET, Central European Time
** Check the different areas & zones on our support site


Outage: Datacenter Sint-Niklaas. SOLVED

Dear customers,

We are currently experiencing an outage with our datacenter in Sint-Niklaas, one of our core routers there has crashed.  Our Network Operations team is investigating this outage.  Further updates will be released as soon as possible.

EDIT 10:19*: The issues were resolved at 9h30. We are investigating further.

Please find all details below:

Start*: 13-08-2021 8:15*
End*: 13-08-2021 9:30*
Duration: Approx. 75 minutes
Impacted services: Mail/webmail services, my edpnet, etc.
Affected Areas**: ALL

 

We apologize for the inconveniences.

Best regards
The edpnet team

*All times are listed in CET, Central European Time
** Check the different areas & zones on our support site